# TimeBlind: A Spatio-Temporal Compositionality Benchmark for Video LLMs

### v1

@@ 1 Introduction
MVBench (li2024mvbench) 4K 2 2 2
TOMATO (shangguan2024tomato) 1.4K 2 4 2 57.3%
Vinoground (zhang2024vinoground) 2K 2 1 2 40.0%
TempCompass (liu2024tempcompass) 4K 2 2 2
@@
Covers all 13 event relations in Allen’s Interval Algebra (allen1983maintaining): before, after, meets, met-by, overlaps, overlapped-by, starts, started-by, finishes, finished-by, during, contains, equals.
@@
Fine-grained spatio-temporal understanding is fundamental for long-horizon video reasoning (wang2025lvbench; pan2025basket) and embodied AI (tang2025video; yang2025cambrian). Beyond simply recognizing what is present, an intelligent system must infer what changes, how it changes, and how multiple changes compose into causal structures. While recent Multimodal Large Language Models (MLLMs) (gemini; openai2025gpt5systemcard; bai2025qwen3vltechnicalreport; clark2025molmo2; cho2025perceptionlm) demonstrate impressive performance on general benchmarks, their “sense of time” remains surprisingly brittle. As shown in Figure 1, even frontier models (e.g., GPT-5 and Gemini 3 Pro) struggle to distinguish atomic actions (e.g., shaking vs. holding) in videos as short as 10 seconds. Moreover, models often misjudge relative dynamics (e.g., accelerating vs. decelerating), and frequently fail to resolve linguistic temporal connectives (e.g., “as soon as”). This discrepancy suggests that current benchmark scores (jang2017tgif; activitynet; li2024mvbench; zhou2025mlvu; fu2025video; hu2025video) severely overestimate genuine temporal understanding capabilities.
@@
The root issue lies in evaluation design. Existing benchmarks (activitynet; li2024mvbench; liu2024tempcompass; fu2025video) rarely isolate temporal structure as the sole discriminative factor. As a result, many models exploit “static shortcuts”—correlating visual entities with answers without modeling time (lei2023revealing; krojer2025shortcut). Furthermore, language priors (li2024naturalbench) allow models to guess answers based on textual plausibility. To truly diagnose understanding, an evaluation must hold the visual content constant and vary only the temporal dynamics.
@@
To address this, we introduce TimeBlind, a diagnostic benchmark for compositional spatio-temporal understanding (Table 1). TimeBlind adopts a minimal-pairs design: each instance contains two videos with near-identical static visual content that differ solely in their temporal structure. To eliminate language priors, we utilize complementary questions where the correct answer flips between the paired videos. By removing static and linguistic shortcuts, TimeBlind forces models to rely exclusively on temporal evidence. Unlike large-scale noisy benchmarks, TimeBlind prioritizes high-fidelity diagnostic precision. Similar to Winoground (thrush2022winoground), each instance serves as a rigorous test for a specific cognitive primitive, where high-quality annotations are prioritized over scale.
@@
Crucially, unlike prior work (zhang2024vinoground; krojer2025shortcut), TimeBlind evaluates logical composition, not just perception. Inspired by cognitive science, we extend the theory of Image Compositionality (krishna2017visual; li2024genai) to the temporal domain, organizing the benchmark around a hierarchical taxonomy: (i) Events, addressing what happened (e.g., changes in object attributes or action understanding); (ii) Event Attributes, describing how events unfold (e.g., speed, magnitude of change); and (iii) Structural Event Logic, examining how multiple events are composed (e.g., temporal topology, causality, and cross-event comparison). Within this framework, we encompass a diverse set of 11 fine-grained categories in real-world scenarios. For example, regarding Temporal Topology in Structural Event Logic, we incorporate Allen’s Interval Algebra (allen1983maintaining) by constructing videos that cover all 13 temporal event relations (e.g., overlaps, meets, equals) and designing questions that distinguish specific topology differences within each pair, going beyond simple sequencing relations (e.g., before, after) used in prior work.
@@
We evaluate over 20 state-of-the-art MLLMs, including leading proprietary models (e.g., Gemini 3 Pro (gemini), GPT-5 (openai2025gpt5systemcard)) as well as strong open-source models (e.g., Qwen3-VL (bai2025qwen3vltechnicalreport), Molmo2 (clark2025molmo2), PLM (cho2025perceptionlm)), on 600 curated instances (totaling 2,400 video-question pairs). We find that even the top-performing model, Gemini-3 Pro, achieves only 48.2% Instance Accuracy (I-Acc), which requires correctly distinguishing both videos in a pair, falling far below human performance (98.2%). Furthermore, while advanced models like Gemini-3 Pro and GPT-5 demonstrate reasonable proficiency in recognizing isolated events (achieving 49.2% and 58.3% I-Acc, respectively), their performance degrades sharply on event attributes—such as Speed (slowly vs. rapidly), and Force (forcefully vs. gently)—dropping to 36.7% and 32.3% I-Acc. Moreover, extensive ablation studies demonstrate that model performance remains poor even with increased input frames or test-time reasoning, as GPT-5 gains only 3.3% I-Acc, indicating models still depend mostly on visual and linguistic shortcuts rather than temporal reasoning. We believe TimeBlind will serve as a vital diagnostic tool for evaluating and developing MLLMs capable of genuine temporal logic.
@@ 2 Related Work
VideoQA Benchmarks. Early VideoQA datasets (activitynet; tgif; perception; xiao2021next; lei2018tvqa; xu2016msr; wu2024star) focus on simple scenarios with short clips and limited question types. Recent benchmarks address comprehensive evaluation (fu2025video; li2024mvbench; ma2025videoeval), complex reasoning (nagrani2025minerva; cheng2025video; hu2025video; song2025video), long-form understanding (mangalam2023egoschema; wang2025lvbench; wu2024longvideobench; song2024moviechat; zhou2025mlvu; yang2025egolife), and domain-specific settings (pan2025basket; long2025adsqa; yi2025exact). However, most do not isolate temporal structure as the sole discriminative factor, allowing models to exploit static shortcuts—relying on object co-occurrence or language priors without genuinely modeling temporal dynamics (lei2023revealing; krojer2025shortcut; li2024naturalbench).
Evaluating Spatio-Temporal Compositionality. To assess the spatio-temporal understanding capabilities of MLLMs, several recent benchmarks have introduced diverse aspects of visual reasoning. In the image understanding domain, BLINK (fu2024blink) reformats and groups classic vision problems into perception-centric multiple-choice questions and carefully removes language priors. In the video understanding domain, several works explicitly target temporal understanding by designing temporally challenging questions (johnson2017clevr; shangguan2024tomato; cai2024temporalbench; xue2025seeing). Beyond single-video evaluation, a growing line of work adopts paired video–question protocols to more precisely diagnose temporal reasoning. TempCompass (liu2024tempcompass) constructs paired videos by systematically manipulating an original video, such as reversing playback or altering temporal speed, thereby isolating temporal understanding from “static shortcuts.” VinoGround (zhang2024vinoground) further strengthens this paradigm by requiring models to answer identical questions over paired videos, where the correct answer is determined solely by temporal differences, effectively removing language priors. Follow-up works like GLIMPSE (zhou2025glimpse) and MVP (krojer2025shortcut) scale this approach to test physical and visual-centric reasoning. Different from prior works, TimeBlind achieves spatio-temporal compositionality through a carefully curated formal taxonomy. Drawing inspiration from cognitive event perception (bach1986algebra) and extending the theory of image compositionality (krishna2017visual; thrush2022winoground; li2024naturalbench), we decompose temporal reasoning into atomic Events, parametric Event Attributes, and Structural Logic.
@@ 3.1 Task Formulation
Diagnostic Metrics. Following rigorous diagnostic protocols (thrush2022winoground; li2024naturalbench), we report a hierarchy of metrics to mitigate shortcuts and assess the models’ genuine temporal understanding:
@@ 3.2 A Hierarchy of Temporal Composition
  • Temporal Topology: We adopt Allen’s Interval Algebra (allen1983maintaining) to comprehensively evaluate interval relations. Unlike prior work limited to simple sequencing (before/after), we cover all 13 relations in our videos, such as overlaps, meets, starts, and during.
@@ 4 Experimental Setup
Baseline Models. To thoroughly assess the challenges posed by TimeBlind, we evaluate over 20 MLLMs, including both frontier proprietary models (e.g., GPT-5 (openai2025gpt5systemcard), Gemini 3 Pro (gemini), Claude (anthropic_claude_sonnet_4_5_2025), and Qwen3-VL Plus (bai2025qwen3vltechnicalreport)) and state-of-the-art open-source models (e.g., Qwen3-VL (bai2025qwen3vltechnicalreport), InternVL 3.5 (internvl3), Molmo2 (clark2025molmo2), PLM (cho2025perceptionlm), Keye-VL (team2025kwai), MiniCPM-V-4.5 (yu2025minicpm), GLM-4.1V (hong2025glm), and Eagle 2.5 (chen2025eagle2.5)).
@@ 5.1 Main Results
Keye-VL-1.5 (team2025kwai) 8B 17.5 26.5 55.6 6.6
LLaVA-Video (zhang2024video) 7B 18.1 34.8 57.8 9.6
InternVL3.5 (internvl3) 8B 30.9 39.0 59.3 13.3
PLM-8B (cho2025perceptionlm) 8B 24.4 32.3 60.6 13.9
Qwen3-VL (bai2025qwen3vltechnicalreport) 8B 34.5 43.6 64.8 19.7
MiniCPM-V-4.5 (yu2025minicpm) 8B 31.7 43.6 62.7 18.3
GLM-4.1V (hong2025glm) 9B 32.3 45.2 62.0 19.7
Qwen3-VL (bai2025qwen3vltechnicalreport) 4B 29.9 43.7 62.6 17.7
VideoChat-Flash (li2024videochat) 7B 31.7 45.4 64.9 19.7
Eagle2.5 (chen2025eagle2.5) 8B 32.2 45.7 64.7 20.2
Molmo2 (clark2025molmo2) 8B 41.0 52.4 68.7 31.2
@@
InternVL 3.5 (internvl3) 30B 31.4 41.3 61.1 17.0
LLaVA-Video (zhang2024video) 72B 28.8 45.0 63.2 18.5
Qwen3-VL (bai2025qwen3vltechnicalreport) 235B 38.1 51.7 66.9 25.8
@@
Claude Sonnet 4.5 (anthropic_claude_sonnet_4_5_2025) - 30.6 38.5 59.2 13.6
Qwen3-VL Plus (bai2025qwen3vltechnicalreport) - 38.7 52.6 67.3 26.0
GPT-5 mini (openai2025gpt5systemcard) - 42.3 53.7 68.3 30.0
Gemini 2.5 Flash (google_gemini_2_5_thinking_2025) - 47.6 53.7 68.9 33.2
Gemini 2.5 Pro (google_gemini_2_5_thinking_2025) - 56.0 62.6 74.2 43.5
GPT-5 (openai2025gpt5systemcard) - 58.5 67.2 77.3 46.3
Gemini 3 Pro (gemini) - 60.2 66.0 76.2 48.2
@@ Appendix A Appendix
This document provides additional details on TimeBlind and failure cases, supplementing the main paper. The outline is as follows:
@@ A.1.1 Data Generation Prompts
In this section, we present the full prompt template used in our Human–AI collaborative pipeline (see Section 3.3 in the main paper).
The system prompt for TimeBlind consists of three parts: (1) a General Prompt that defines the role and task, (2) Detailed Requirements that specify the core generation rules and logical constraints, and (3) an Output Format that standardizes the response structure. The prompt includes strict structural constraints to ensure high-quality minimal pairs. When using this prompt, we provide a single category as the “Input” each time.
@@ A.1.3 Taxonomy
Table 8: Skill Definitions in TimeBlind for Video Event Understanding.

### v2

@@ 1 Introduction
MVBench (Li et al., 2024c) 4K 2 2 2
TOMATO (Shangguan et al., 2024) 1.4K 2 4 2 57.3%
Vinoground (Zhang et al., 2024a) 2K 2 1 2 40.0%
TempCompass (Liu et al., 2024) 4K 2 2 2
@@
Covers all 13 event relations in Allen’s Interval Algebra (Allen, 1983): before, after, meets, met-by, overlaps, overlapped-by, starts, started-by, finishes, finished-by, during, contains, equals.
@@
Fine-grained spatio-temporal understanding is fundamental for long-horizon video reasoning (Wang et al., 2025b; Pan et al., 2025) and embodied AI (Tang et al., 2025; Yang et al., 2025b). Beyond simply recognizing what is present, an intelligent system must infer what changes, how it changes, and how multiple changes compose into causal structures. While recent Multimodal Large Language Models (MLLMs) (Pichai et al., 2025; OpenAI, 2025; Bai et al., 2025; Clark et al., 2025; Cho et al., 2025) demonstrate impressive performance on general benchmarks, their “sense of time” remains surprisingly brittle. As shown in Figure 1, even frontier models (e.g., GPT-5 and Gemini 3 Pro) struggle to distinguish atomic actions (e.g., shaking vs. holding) in videos as short as 10 seconds. Moreover, models often misjudge relative dynamics (e.g., accelerating vs. decelerating), and frequently fail to resolve linguistic temporal connectives (e.g., “as soon as”). This discrepancy suggests that current benchmark scores (Jang et al., 2017a; Yu et al., 2019; Li et al., 2024c; Zhou et al., 2025a; Fu et al., 2025; Hu et al., 2025) severely overestimate genuine temporal understanding capabilities.
@@
The root issue lies in evaluation design. Existing benchmarks (Yu et al., 2019; Li et al., 2024c; Liu et al., 2024; Fu et al., 2025) rarely isolate temporal structure as the sole discriminative factor. As a result, many models exploit “static shortcuts”—correlating visual entities with answers without modeling time (Lei et al., 2023; Krojer et al., 2025). Furthermore, language priors (Li et al., 2024b) allow models to guess answers based on textual plausibility. To truly diagnose understanding, an evaluation must hold the visual content constant and vary only the temporal dynamics.
@@
To address this, we introduce TimeBlind, a diagnostic benchmark for compositional spatio-temporal understanding (Table 1). TimeBlind adopts a minimal-pairs design: each instance contains two videos with near-identical static visual content that differ solely in their temporal structure. To eliminate language priors, we utilize complementary questions where the correct answer flips between the paired videos. By removing static and linguistic shortcuts, TimeBlind forces models to rely exclusively on temporal evidence. Unlike large-scale noisy benchmarks, TimeBlind prioritizes high-fidelity diagnostic precision. Similar to Winoground (Thrush et al., 2022), each instance serves as a rigorous test for a specific cognitive primitive, where high-quality annotations are prioritized over scale.
@@
Crucially, unlike prior work (Zhang et al., 2024a; Krojer et al., 2025), TimeBlind evaluates logical composition, not just perception. Inspired by cognitive science, we extend the theory of Image Compositionality (Krishna et al., 2017; Li et al., 2024a) to the temporal domain, organizing the benchmark around a hierarchical taxonomy: (i) Events, addressing what happened (e.g., changes in object attributes or action understanding); (ii) Event Attributes, describing how events unfold (e.g., speed, magnitude of change); and (iii) Structural Event Logic, examining how multiple events are composed (e.g., temporal topology, causality, and cross-event comparison). Within this framework, we encompass a diverse set of 11 fine-grained categories in real-world scenarios. For example, regarding Temporal Topology in Structural Event Logic, we incorporate Allen’s Interval Algebra (Allen, 1983) by constructing videos that cover all 13 temporal event relations (e.g., overlaps, meets, equals) and designing questions that distinguish specific topology differences within each pair, going beyond simple sequencing relations (e.g., before, after) used in prior work.
@@
We evaluate over 20 state-of-the-art MLLMs, including leading proprietary models (e.g., Gemini 3 Pro (Pichai et al., 2025), GPT-5 (OpenAI, 2025)) as well as strong open-source models (e.g., Qwen3-VL (Bai et al., 2025), Molmo2 (Clark et al., 2025), PLM (Cho et al., 2025)), on 600 curated instances (totaling 2,400 video-question pairs). We find that even the top-performing model, Gemini-3 Pro, achieves only 48.2% Instance Accuracy (I-Acc), which requires correctly distinguishing both videos in a pair, falling far below human performance (98.2%). Furthermore, while advanced models like Gemini-3 Pro and GPT-5 demonstrate reasonable proficiency in recognizing isolated events (achieving 49.2% and 58.3% I-Acc, respectively), their performance degrades sharply on event attributes—such as Speed (slowly vs. rapidly), and Force (forcefully vs. gently)—dropping to 36.7% and 32.3% I-Acc. Moreover, extensive ablation studies demonstrate that model performance remains poor even with increased input frames or test-time reasoning, as GPT-5 gains only 3.3% I-Acc, indicating models still depend mostly on visual and linguistic shortcuts rather than temporal reasoning. We believe TimeBlind will serve as a vital diagnostic tool for evaluating and developing MLLMs capable of genuine temporal logic.
@@ 2 Related Work
VideoQA Benchmarks. Early VideoQA datasets (Yu et al., 2019; Jang et al., 2017b; Patraucean et al., 2023; Xiao et al., 2021; Lei et al., 2018; Xu et al., 2016; Wu et al., 2024a) focus on simple scenarios with short clips and limited question types. Recent benchmarks address comprehensive evaluation (Fu et al., 2025; Li et al., 2024c; Ma et al., 2025), complex reasoning (Nagrani et al., 2025; Cheng et al., 2025; Hu et al., 2025; Song et al., 2025), long-form understanding (Mangalam et al., 2023; Wang et al., 2025b; Wu et al., 2024b; Song et al., 2024; Zhou et al., 2025a; Yang et al., 2025a), and domain-specific settings (Pan et al., 2025; Long et al., 2025; Yi et al., 2025). However, most do not isolate temporal structure as the sole discriminative factor, allowing models to exploit static shortcuts—relying on object co-occurrence or language priors without genuinely modeling temporal dynamics (Lei et al., 2023; Krojer et al., 2025; Li et al., 2024b).
Evaluating Spatio-Temporal Compositionality. To assess the spatio-temporal understanding capabilities of MLLMs, several recent benchmarks have introduced diverse aspects of visual reasoning. In the image understanding domain, BLINK (Fu et al., 2024) reformats and groups classic vision problems into perception-centric multiple-choice questions and carefully removes language priors. In the video understanding domain, several works explicitly target temporal understanding by designing temporally challenging questions (Johnson et al., 2017; Shangguan et al., 2024; Cai et al., 2024; Xue et al., 2025). Beyond single-video evaluation, a growing line of work adopts paired video–question protocols to more precisely diagnose temporal reasoning. TempCompass (Liu et al., 2024) constructs paired videos by systematically manipulating an original video, such as reversing playback or altering temporal speed, thereby isolating temporal understanding from “static shortcuts.” VinoGround (Zhang et al., 2024a) further strengthens this paradigm by requiring models to answer identical questions over paired videos, where the correct answer is determined solely by temporal differences, effectively removing language priors. Follow-up works like GLIMPSE (Zhou et al., 2025b) and MVP (Krojer et al., 2025) scale this approach to test physical and visual-centric reasoning. Different from prior works, TimeBlind achieves spatio-temporal compositionality through a carefully curated formal taxonomy. Drawing inspiration from cognitive event perception (Bach, 1986) and extending the theory of image compositionality (Krishna et al., 2017; Thrush et al., 2022; Li et al., 2024b), we decompose temporal reasoning into atomic Events, parametric Event Attributes, and Structural Logic.
@@ 3.1 Task Formulation
Diagnostic Metrics. Following rigorous diagnostic protocols (Thrush et al., 2022; Li et al., 2024b), we report a hierarchy of metrics to mitigate shortcuts and assess the models’ genuine temporal understanding:
@@ 3.2 A Hierarchy of Temporal Composition
  • Temporal Topology: We adopt Allen’s Interval Algebra (Allen, 1983) to comprehensively evaluate interval relations. Unlike prior work limited to simple sequencing (before/after), we cover all 13 relations in our videos, such as overlaps, meets, starts, and during.
@@ 4 Experimental Setup
Baseline Models. To thoroughly assess the challenges posed by TimeBlind, we evaluate over 20 MLLMs, including both frontier proprietary models (e.g., GPT-5 (OpenAI, 2025), Gemini 3 Pro (Pichai et al., 2025), Claude (Anthropic, 2025), and Qwen3-VL Plus (Bai et al., 2025)) and state-of-the-art open-source models (e.g., Qwen3-VL (Bai et al., 2025), InternVL 3.5 (Wang et al., 2025a), Molmo2 (Clark et al., 2025), PLM (Cho et al., 2025), Keye-VL (Team et al., 2025), MiniCPM-V-4.5 (Yu et al., 2025), GLM-4.1V (Hong et al., 2025), and Eagle 2.5 (Chen et al., 2025)).
@@ 5.1 Main Results
Keye-VL-1.5 (Team et al., 2025) 8B 17.5 26.5 55.6 6.6
LLaVA-Video (Zhang et al., 2024b) 7B 18.1 34.8 57.8 9.6
InternVL3.5 (Wang et al., 2025a) 8B 30.9 39.0 59.3 13.3
PLM-8B (Cho et al., 2025) 8B 24.4 32.3 60.6 13.9
Qwen3-VL (Bai et al., 2025) 8B 34.5 43.6 64.8 19.7
MiniCPM-V-4.5 (Yu et al., 2025) 8B 31.7 43.6 62.7 18.3
GLM-4.1V (Hong et al., 2025) 9B 32.3 45.2 62.0 19.7
Qwen3-VL (Bai et al., 2025) 4B 29.9 43.7 62.6 17.7
VideoChat-Flash (Li et al., 2024d) 7B 31.7 45.4 64.9 19.7
Eagle2.5 (Chen et al., 2025) 8B 32.2 45.7 64.7 20.2
Molmo2 (Clark et al., 2025) 8B 41.0 52.4 68.7 31.2
@@
InternVL 3.5 (Wang et al., 2025a) 30B 31.4 41.3 61.1 17.0
LLaVA-Video (Zhang et al., 2024b) 72B 28.8 45.0 63.2 18.5
Qwen3-VL (Bai et al., 2025) 235B 38.1 51.7 66.9 25.8
@@
Claude Sonnet 4.5 (Anthropic, 2025) - 30.6 38.5 59.2 13.6
Qwen3-VL Plus (Bai et al., 2025) - 38.7 52.6 67.3 26.0
GPT-5 mini (OpenAI, 2025) - 42.3 53.7 68.3 30.0
Gemini 2.5 Flash (Kavukcuoglu, 2025) - 47.6 53.7 68.9 33.2
Gemini 2.5 Pro (Kavukcuoglu, 2025) - 56.0 62.6 74.2 43.5
GPT-5 (OpenAI, 2025) - 58.5 67.2 77.3 46.3
Gemini 3 Pro (Pichai et al., 2025) - 60.2 66.0 76.2 48.2
@@ Appendix A Appendix
This document details TimeBlind and failure cases, supplementing the main paper. The outline is as follows:
@@ A.1.1 Data Generation Prompts
In this section, we present the full prompt template used in our Human–AI collaborative pipeline (see Section 3.3). The system prompt for TimeBlind consists of three parts: (1) a General Prompt that defines the role and task, (2) Detailed Requirements that specify the core generation rules and logical constraints, and (3) an Output Format that standardizes the response structure. The prompt includes strict structural constraints to ensure high-quality minimal pairs. When using this prompt, we provide a single category as the “Input” each time.
@@ A.1.3 Taxonomy
Table 8: Hierarchical taxonomy of TimeBlind for video understanding. The taxonomy is structured into three levels—Event, Event Attribute, and Structural Event Logic. For each level, we present its conceptual definition, associated sub-categories, and representative paired-video descriptions for each sub-category.